Billie combined 200 grams of N2 and 100 grams H2 to produce 230 grams of ammonia. Determines the limiting reagent, excess reagent, percent yield and theoretical yield of the reaction. show the steps on how you came up with your answer


1
Expert's answer
2022-04-14T01:17:06-0400

N2 + 3H2 = 2NH3

n(N2)=m(N2)/Mr(N2)=200g / 28g/moll = 7.14 moles

n(H2) = m(H2)/3Mr(H2) = 100g / 6g/moll = 16.67 moles

H2 excess reagent

Nlimiting reagent


m(NH3= ( m(N2)*Mr(NH3) ) / Mr(N2) = (200g * 34g/moll) / 28g/milk = 242.86g - theoretical yield


percent yield = (230/242.86)*100% = 94.7%